OpenStudy (anonymous):

9=sqrt(3)^(4x+6) solve

OpenStudy (anonymous):

(4x+6)ln(sqrt(3))=ln(9) 4x+6=ln(9)/ln(sqrt(3)) 4x=ln(9)/ln(sqrt(3))-6 x=ln(9)/[4*ln(sqrt(3))]-3/2